OpenStudy (anonymous):

Suppose that \(p\) is a prime and \( p>3 \).Prove that each divisors of \(\large \frac{ 2^p+1}{ 3 }\) can be written in the form of \( 2pk + 1\) (\(k\) is a natural number).

OpenStudy (anonymous):

for example if \(p=5\) then \(11 = 2(5) + 1\).

OpenStudy (shubhamsrg):

is it 2px + 1 , or 2pk + 1?

OpenStudy (anonymous):

Oops ! *edited

OpenStudy (shubhamsrg):

We can try it by induction, hope induction is allowed ?

OpenStudy (shubhamsrg):

well nop, i don't think that is a good idea, because prime numbers are not continuosly distributed..sorry

OpenStudy (shubhamsrg):

2^p +1 = (3-1)^p -1 = -1 + C(p,1) 3 - C(p,2) 3^2 .... + 3^p +1 (-1)^p = -1 because p is odd (obviously) => 2^p +1 = C(p,1) 3 - C(p,2) 3^2 .... C(p,p) 3^p (2^p +1)/3 = C(p,1) - C(p,2) 3 .... 3^(p-1)

OpenStudy (anonymous):

hmm,why do u try hard ways ? :D

ganeshie8 (ganeshie8):

only prime divisors are of form 2pk+1 right ?

ganeshie8 (ganeshie8):

actually if prime divisors are of form 2pk+1, all other divisors also will be of form 2pk+1 i guess ?

OpenStudy (anonymous):

yes ;)

ganeshie8 (ganeshie8):

then it is sufficient to prove it for prime divisors

OpenStudy (anonymous):

right ;)

OpenStudy (shubhamsrg):

let q = C(p,1) - C(p,2) 3 .... 3^(p-1) q = p j + 3^(p-1) where j is some natural number. Now we need to prove two things, 3^(p-1) -1 is always divisible by p. and q is always even. hmm

OpenStudy (shubhamsrg):

not, q , but q-1

ganeshie8 (ganeshie8):

let \(q \gt 3\) is a prime divisor of \(\frac{2^p+1}{3}\) \(\large q | \frac{2^p+1}{3}\) and \(q \gt 3\) together imply that \(q | 2^p+1\) yes ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

thats almost same as what shubham arrived at..

ganeshie8 (ganeshie8):

so we have \[2^p+1\equiv 0 \pmod{q}\] which is same as \[2^p\equiv -1 \pmod{q}\]

ganeshie8 (ganeshie8):

that means the order of \(2\) in modulo \(q\) is \(2p\) By definition the order divides \(\phi(q)\) : \[2p|\phi(q)\\2p|q-1\\q = 2pk+1\] for some integer \(k\)

OpenStudy (anonymous):

nice ;) i just expected that proof ;)

ganeshie8 (ganeshie8):

you need to work \(q=3\) case separately as we have assumed \(q\gt 3\) above

OpenStudy (anonymous):

but u only said the order of 2 in modulo q = 2p,i don't know how did u prove that but my proof is that if we suppose the order of 2 in mudulo q = r then we can say \(r|2p \)

OpenStudy (anonymous):

we have 4 cases: \(r=1\), \(r=p\) ,\(r=2\) , \(r=2p\)

OpenStudy (anonymous):

we can say why r cannot be 1,p,2.so it's 2p.

OpenStudy (anonymous):

for r=1 we have q|1,for r=2 we have q|3 so q|3 we can say why it cannot happen.and for r=p we have \(q|2^p - 1 \) and \(q|2^p +1 \) so \( q|2\)

ganeshie8 (ganeshie8):

that looks neat xD

OpenStudy (anonymous):

if \(q=3\) then \( 2^{2p} \equiv 1 ~ (mod 9) \) , but the order of 2 mod 9 is 6,so \(6|2p\) then \(3|p\) but p cannot be 3.

ganeshie8 (ganeshie8):

\[2^p \equiv -1 \pmod{q}\] another way to argue the order cannot be \(\le p\) is by noticing the "-1" on right hand side if the order were 1 or p, then any power of it yields "1". But we have \(2^p \not \equiv 1 \pmod{q}\). This precicely allows us to conclude that the order of \(2\) cannot be \(\le p\)

OpenStudy (anonymous):

and \( p>3 \) so the order is greater than 3! yes ,nice idea ;)

ganeshie8 (ganeshie8):

we're assuming \(q\gt 3\)